OpenStudy (anonymous):

My question is: The following table of the years of experience of university teachers and their corresponding salary in thousands of dollars. Years of experience /salary 0yrs 58,000 5 yrs 68,000 10 yrs 78,000 a. Give the linear function S(t) that gives the annual salary in thousands of dollars for t years of experience b. predict the average salary for a teacher with 15 years experience ( i think it is 82,000) but don't know how to work the problem c. Find the years of experience for a teacher that is paid 84,000. (I think 12.5 but don't know how to show the work)

OpenStudy (anonymous):

So, you will need to find the linear equation using the points given to you. What points? Well, lets see ;) 0 yrs, 58,000....This looks like a point! (0, 58 000). 5 yrs, 68,000....Another point! (5, 68000) First find the gradient: \[m = \frac{y _{2} - y_{1}}{x_{2}-x_{1}}\] Then find the equation of the line using the formula\[y -y_{1} = m(x-x_{1})\] If you need help solving the equations or using the formulas post back :)

OpenStudy (anonymous):

Alright, so as I said we first need to find the gradient. We have our points to do that. (0,58000) and (5,68000). (x1, y1) (x2,y2) Substitute these into the formula \[m = \frac{68000 - 58000}{5-0}\]\[m = 2,000\] Now we must use the linear equation formula: \[y - 58,000 = 2,000(x-0)\]\[y = 2,000x + 58,000\] Change 'y' to S(t) and 'x' to t.\[S(t) = 2,000(t) + 58,000\] For question b: Substitute the time, 15 years, into the formula (so t = 15). \[S(15) = 2,000(15) + 58,000\]\[S(15) = 88,000\] For question c: Change S(t) to 84,000 and solve. \[84,000 = 2,000t + 58000\]\[84,000 - 58,000 = 2,000t\]\[26,000 = 2,000t\]\[26,000 \div 2,000 = t\]\[t = 13\] So, the answer is 13 years for c.

OpenStudy (anonymous):

Thank you.